OpenStudy (adll23):

The perimeter of a rectangle is 64 units. Can the length x of the rectangle can be 20 units when its width y is 11 units? (4 points) No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 64 No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 32 Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 64 Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 32

OpenStudy (narutoboy14):

d

OpenStudy (mathstudent55):

What is the perimeter of a rectangle?

OpenStudy (adll23):

2 (w + h)

OpenStudy (mathstudent55):

Great. The problem is calling the length x and the width y. According to the variables of the problem, then, the perimeter is P = 2(x + y) Ok?

OpenStudy (mathstudent55):

We are told the length is 20 and the width is 11. What would the perimeter be? P = 2(x + y) P = 2(20 + 11) P = 2(31) P = 62 With the length 20 and width 11, the perimeter is 62, not 64. This means the perimeter cannot be 64.

OpenStudy (mathstudent55):

Notice that when you calculate the perimeter, you add the length and width, and then you multiply that by 2. That means the length + the width must equal half the perimeter. In our case, the length + the width = 20 + 11 = 31 Since we are told the perimeter is 64, half the perimeter is 32. 31 is not half the perimeter, so 64 cannot be the perimeter.

OpenStudy (adll23):

So it would be B?

OpenStudy (adll23):

@mathstudent55

OpenStudy (adll23):

Or would it be A? @mathstudent55

OpenStudy (narutoboy14):

b

OpenStudy (mathstudent55):

The answer is B.